# Quantum Non-Local Magic

### v2

@@ Abstract
We then consider the MHC at a critical measurement rate of $p=0.17$. The circuit structure is shown in Fig. 3(a). We consider the average NM defined as ¯$\overline{\mathcal{M}}=\mathbb{E}_{U}\mathbb{E}_{m}[\mathcal{M}]$, where $\mathbb{E}_{U}$ and $\mathbb{E}_{m}$ represent the averaging over unitary gates and measurement outcomes, respectively. Since the model is structureless, with only locality as a constraint, $\rho_{1,r}$ does not necessarily have a canonical Pauli basis representation, and the NM must be computed by numerical optimization. As shown in Fig. 3(b), NM is significantly lower than the total magic and also feature a power-law decay. We then consider the MINM and compare it with the average total magic for post-measurement states, as presented in Fig. 3(c). The measurement direction is inessential, and we assume it to be along the $z$-axis. It is evident that a greater amount of magic can be removed by local unitary transformations as the distance increases. Furthermore, we observe that the decay exponent of MINM is smaller than half of that of mutual information prior to measurement, as shown in Fig. 3(d). Since the slowest correlation a state can maintain must decay at an exponent at least half that of mutual information [70], this suggests that the measurement induces NM with a longer range than any pre-existing correlations. This phenomenon is akin to entanglement swapping, where measurements induce long-range entanglement between previously uncorrelated qubits. We refer to this effect as “magic swapping”. Since it has been established that $\mathcal{M}$ is upper-bounded by entanglement [35], and that measurement-induced entanglement cannot decay more slowly than correlations in sign-free states [71], it follows that magic swapping can only occur in states possessing sign structures.
@@
Discussions. As we focus on the NM for qubit systems, a natural extension is to investigate systems with higher-dimensional local Hilbert spaces, such as qutrits. In systems where the local Hilbert space dimension is an odd prime, mana serves as an alternative, appropriate magic measure for both pure and mixed states [72, 73, 74, 75]. Notably, like SRE, mana does not necessitate any minimization procedure. However, it remains an open question whether an analytical expression can be derived for the two-qutrit case. It would also be interesting to investigate the consistency between the results obtained using SRE and mana. We leave this to future work.
@@
While NM can be expressed analytically for two-qubit pure states, making it computationally straightforward and feasible for experimental measurements, the situation becomes more complex for general multi-qubit states. In this case, computing the NM presents an optimization problem involving $O(N^{2})$ parameters, which becomes increasingly difficult as the number of qubits grows. One approach to address this challenge is to employ automatic differentiation techniques, similar to those used in simulating the variational quantum eigensolver algorithm, with the target function being the total magic rather than the energy [76]. Scaling up the calculation of NM would facilitate the classification of states based on whether they exhibit volume-law or area-law NM, analogous to the classification of states in terms of entanglement. We conjecture that only states that exhibit both volume-law entanglement entropy and volume-law magic can sustain a volume-law NM.
@@
Additionally, it is noteworthy that a related measure, non-stabilizerness entanglement entropy (NsEE), has recently been introduced [77, 78]. In this formulation, free operations are defined as global Clifford gates instead of local unitary gates, and the quantity calculated is the entanglement entropy of the total state, rather than its magic. While NsEE and NM aim to characterize the “quantumness” of a system from distinct perspectives, further investigation into the relationship between these two measures would be a valuable direction for future research.
@@ Acknowledgements.
Supplemental Material for “Quantum Non-Local Magic”
@@
S1.1 Proof of subadditivity
In this section, we prove that the non-local magic (NM), defined as $\mathcal{M}_{AB}(\left|\Psi_{AB}\right\rangle)=\min_{U_{A}\otimes U_{B}}M(U_{A}\otimes U_{B}\left|\Psi_{AB}\right\rangle)$, is subadditive if the magic measure $M$ is additive. By definition, we have:
@@
This establishes the subadditivity of $\mathcal{M}$.
Building on this subadditivity, it is straightforward to establish an upper bound for the NM of a many-body state, provided that it can be brought into product states of two-qubit pure states in canonical form via local unitary transformations. It has been shown that such transformations can be approximately realized using the Petz recovery map for Haar-random states [1]. However, the exact determination of such a transformation remains an open problem and is a subject worth further exploration.
@@
where $S_{2}(\rho)$ denotes the 2-Rényi entropy of the state $\rho$, which is invariant under any unitary transformation applied to $\rho$. Therefore, the calculation of NM $\mathcal{M}$ reduces to finding the appropriate local unitary transformation that maximizes the sum $\sum_{\mu\nu}{|R_{\mu\nu}|^{4}}$.
@@
where the final equality follows from the orthogonality of $O_{A}$. A similar argument holds for the action of $O_{B}$. This indicates that a diagonal matrix $T$ maximizes the fourth power sum of its elements under any local unitary transformation. Combining these results, we conclude that any local unitary transformation applied to a two-qubit state with fully polarized Bloch vectors for both qubits and a diagonal correlation matrix increases its magic. We call such a state to be in the canonical form of its Pauli basis representation. It follows directly that if a state $\rho$ is in canonical form, then $\mathcal{M}(\rho)=M(\rho)$.
@@
In this section, we derive the analytic expression for the probability distribution of NM for the two-qubit Haar random state. It is well known that the Haar random state has an entanglement spectrum $\{\text{cos}(\theta),\text{sin}(\theta)\}$, with the associated probability distribution given by $P(x)=3(1-2x)^{2}$, where we define $x\equiv\text{cos}^{2}(\theta)$ [5, 6]. From the corollary stated in the main text, the NM for a state with such an entanglement spectrum is expressed as:
@@
In this section, we demonstrate that the decay exponent of NM in the critical ground state of TFIM is identical to that of mutual information. We begin by deriving a general result for the decay exponent of the two-point total magic. At the critical point, all connected correlation functions of the form $\mathcal{C}_{\mu\nu}=\text{Tr}(\rho_{1,r}\sigma_{\mu}^{1}\sigma_{\nu}^{r})-\text{Tr}(\rho_{1}\sigma_{\mu}^{1})\text{Tr}(\rho_{r}\sigma_{\nu}^{r})$ obey a power-law decay, i.e., $\mathcal{C}_{\mu\nu}\sim r^{-\alpha_{\mu\nu}}$, where $\mu,\nu$ takes values in $\{0,x,y,z\}$. For the correlations that vanish, we assume that the corresponding decay exponent $\alpha$ is infinite. Thus, we can write the correlation function as:
@@
where $c_{\mu\nu}\equiv\text{Tr}(\rho_{1}\sigma_{\mu}^{1})\text{Tr}(\rho_{r}\sigma_{\nu}^{r})$ is independent of $r$ due to translational invariance. Using this, the total magic of the state can be expressed as:
@@
where $(1+x)^{-1}\sim(1-x)$ and $\text{log}(1+x)\sim x$ for small $x$ are used. Denoting $\alpha^{*}\equiv\min_{\mu\nu}\alpha_{\mu\nu}$ and $\overline{\alpha}\equiv\min_{\mu\nu}\{\alpha_{\mu\nu};c_{\mu\nu}\neq 0\}$, it follows that the in the large $r$ limit, the total magic $M$ is governed by:
@@
Thus, the decay exponent of the total magic is given by:
@@
At the critical point of the TFIM, we have $\alpha^{*}=\alpha_{xx}=1/4$ and $\overline{\alpha}=\alpha_{zz}=2$ [7], which results in $\alpha_{M}=1/2$, identical to the decay exponent of the mutual information. Furthermore, since the ground state is in canonical form, the NM is equivalent to the total magic of the two-qubit state. This completes the proof.
@@
In this section, we provide numerical details for the simulations conducted in the main text. The ground state of the TFIM is obtained using the density matrix renormalization group algorithm using the ITensor package [8, 9], with the maximum bond dimension $\chi=500$ and truncation cutoff $\epsilon=10^{-12}$. For the monitored Haar-random unitary circuit at the critical point, the time evolution was performed by the time-evolving block decimation algorithm, with the maximum bond dimension $\chi=1000$ and truncation cutoff $\epsilon=10^{-6}$.
When calculating NM through numerical optimization over all local unitary transformations, we parameterize the local unitary on a single qubit as $U=\begin{pmatrix}\text{cos}(\theta/2)&-e^{i\lambda}\text{sin}(\theta/2)\\ e^{i\phi}\text{sin}(\theta/2)&e^{i(\phi+\lambda)}\text{cos}(\theta/2)\end{pmatrix}$, where $\phi\in[0,2\pi]$, $\theta\in[0,\pi]$ and $\lambda\in[0,2\pi]$. The optimization process employed the Nelder-Mead algorithm, as implemented in the Optim package [10]. To mitigate the risk of converging to a local minimum, we initiated the optimization from 100 distinct starting points and selected the minimal value achieved.
S2.2 Separable mixed states host non-zero NM
In this section, we demonstrate that for any choice of magic measure for mixed states, certain separable states will always exhibit a non-zero NM. A separable state is defined as the convex hull of pure product states [11]:
@@
where $\psi_{i}$ are pure stabilizer states. When STAB is regarded as the set of free states, it specifically captures distillable non-Clifford resources. On the other hand, regarding $\text{STAB}_{0}$ as free states, it also incorporates the non-flat probability distributions of mixed states as resources. In the second SRE used in the main text, $\text{STAB}_{0}$ is treated as the set of free states, while other magic measures, such as the robustness of magic (RoM) [14, 15, 16] and mana [17], consider STAB as the set of free states. It is also worth mentioning that the SRE has been extended by a convex-roof construction to provide an alternative measure in the latter case [18].
@@
This scenario changes when STAB is considered as the set of free states, as any single-qubit density matrix can always be diagonalized, thus ensuring that it lies within STAB. However, we are unaware of any proof guaranteeing that every separable multi-qubit state can be transformed into a state belonging to STAB using only local unitary transformations, and we conjecture that this is not the case. To provide numerical evidence, consider the separable state:
@@
Here, $\left|\phi_{0}\right\rangle=\text{cos}(\pi/8)\left|0\right\rangle+\text{sin}(\pi/8)\left|1\right\rangle$. Using RoM as the magic measure, which is defined as $R(\rho)=\min_{x}\{\sum_{i}|x_{i}|;\rho=\sum_{i}x_{i}\left|\psi_{i}\right\rangle\left\langle\psi_{i}\right|\}-1$, we numerically minimize the magic over all possible local unitary transformations and find that $\mathcal{M}_{R}(\rho_{0})\approx 0.0703$, which is non-zero. This indicates that even when using magic measures that treat STAB as the set of free states, certain separable states can still exhibit a non-zero NM.

### v3

@@ Abstract
We then consider the MHC at a critical measurement rate of $p=0.17$. The circuit structure is shown in Fig. 3(a). We consider the average NN defined as ¯$\overline{\mathcal{M}}=\mathbb{E}_{U}\mathbb{E}_{m}[\mathcal{M}]$, where $\mathbb{E}_{U}$ and $\mathbb{E}_{m}$ represent the averaging over unitary gates and the measurement outcomes, respectively. Since the model is structureless, with only locality as a constraint, $\rho_{1,r}$ does not necessarily have a canonical Pauli basis representation, and the NN must be computed by numerical optimization. As shown in Fig. 3(b), NN is significantly lower than total nonstabilizerness and also exhibits a power-law decay. We then consider the MINN and compare it with the average total nonstabilizerness for post-measurement states, as presented in Fig. 3(c). The measurement direction is insignificant, and we assume it to be along the $z$-axis. It is evident that a greater amount of nonstabilizerness can be removed by local unitary transformations as the distance increases. Furthermore, we observe that the decay exponent of MINN is less than half of that of mutual information prior to measurement, as shown in Fig. 3(d). Since the slowest correlation a state can maintain must decay at an exponent at least half that of mutual information [74], this suggests that the measurement induces NN with a longer range than any pre-existing correlations. This phenomenon is akin to entanglement swapping, where measurements induce long-range entanglement between previously uncorrelated qubits. We refer to this effect as “nonstabilizerness swapping”. Since it has been established that $\mathcal{M}$ is upper bounded by entanglement [38], and that measurement-induced entanglement cannot decay more slowly than correlations in sign-free states [75], it follows that nonstabilizerness swapping can only occur in states possessing sign structures.
@@
Discussions. As we focus on the NN for qubit systems, a natural extension is to investigate systems with higher-dimensional local Hilbert spaces, such as qutrits. In systems where the local Hilbert space dimension is an odd prime, mana serves as an alternative, appropriate nonstabilizerness measure for both pure and mixed states [76, 77, 78, 79]. Notably, like SRE, mana does not necessitate any minimization procedure. However, it remains an open question whether an analytical expression can be derived for the two-qutrit case. It would also be interesting to investigate the consistency between the results obtained using SRE and mana. We leave this to future work.
@@
While NN can be expressed analytically for two-qubit pure states, making it computationally straightforward and feasible for experimental measurements, the situation becomes more complex for general multi-qubit states. In this case, computing the NN presents an optimization problem involving $O(N^{2})$ parameters, which becomes increasingly difficult as the number of qubits grows. An approach to address this challenge is to employ automatic differentiation techniques, similar to those used in simulating the variational quantum eigensolver algorithm, with the target function being the total nonstabilizerness rather than the energy [80]. Scaling up the calculation of NN would facilitate the classification of states based on whether they exhibit volume-law or area-law NN, analogous to the classification of states in terms of entanglement. We conjecture that only states that exhibit both volume-law entanglement entropy and volume-law nonstabilizerness can sustain a volume-law NN.
@@
Additionally, it is noteworthy that a related measure, nonstabilizerness entanglement entropy (NsEE), has recently been introduced [81, 82]. In this formulation, free operations are defined as global Clifford gates instead of local unitary gates, and the computed quantity is the entanglement entropy of the total state rather than its nonstabilizerness. While NsEE and NN aim to characterize the “quantumness” of a system from distinct perspectives, further investigation into the relationship between these two measures would be a valuable direction for future research.
@@ Acknowledgements.
Supplemental Material for “Quantum Non-Local Nonstabilizerness”
@@
S1.1 Proof of sub-additivity
In this section, we prove that the non-local nonstabilizerness (NN), defined as $\mathcal{M}_{AB}(\left|\Psi_{AB}\right\rangle)=\min_{U_{A}\otimes U_{B}}M(U_{A}\otimes U_{B}\left|\Psi_{AB}\right\rangle)$, is sub-additive if the nonstabilizerness measure $M$ is additive. By definition, we have:
@@
This establishes the sub-additivity of $\mathcal{M}$.
Building on this sub-additivity, it is straightforward to establish an upper bound for the NN of a many-body state, provided that it can be brought into product states of two-qubit pure states in canonical form via local unitary transformations. It has been shown that such transformations can be approximately realized using the Petz recovery map for Haar-random states [1]. However, the exact determination of such a transformation remains an open problem and is a subject worth further exploration.
@@
where $S_{2}(\rho)$ denotes the 2-Rényi entropy of the state $\rho$, which is invariant under any unitary transformation applied to $\rho$. Therefore, the calculation of NN $\mathcal{M}$ reduces to finding the appropriate local unitary transformation that maximizes the sum $\sum_{\mu\nu}{|R_{\mu\nu}|^{4}}$.
@@
where the final equality follows from the orthogonality of $O_{A}$. A similar argument holds for the action of $O_{B}$. This indicates that a diagonal matrix $T$ maximizes the fourth power sum of its elements under any local unitary transformation. Combining these results, we conclude that any local unitary transformation applied to a two-qubit state with fully polarized Bloch vectors for both qubits and a diagonal correlation matrix increases its nonstabilizerness. We call such a state to be in the canonical form of its Pauli basis representation. It follows directly that if a state $\rho$ is in canonical form, then $\mathcal{M}(\rho)=M(\rho)$.
@@
In this section, we derive the analytic expression for the probability distribution of NN for the two-qubit Haar random state. It is well known that the Haar random state has an entanglement spectrum $\{\text{cos}(\theta),\text{sin}(\theta)\}$, with the associated probability distribution given by $P(x)=3(1-2x)^{2}$, where we define $x\equiv\text{cos}^{2}(\theta)$ [5, 6]. From the corollary stated in the main text, the NN for a state with such an entanglement spectrum is expressed as:
@@
In this section, we demonstrate that the decay exponent of NN in the critical ground state of TFIM is identical to that of mutual information. We begin by deriving a general result for the decay exponent of the two-point total nonstabilizerness. At the critical point, all connected correlation functions of the form $\mathcal{C}_{\mu\nu}=\text{Tr}(\rho_{1,r}\sigma_{\mu}^{1}\sigma_{\nu}^{r})-\text{Tr}(\rho_{1}\sigma_{\mu}^{1})\text{Tr}(\rho_{r}\sigma_{\nu}^{r})$ obey a power-law decay, i.e., $\mathcal{C}_{\mu\nu}\sim r^{-\alpha_{\mu\nu}}$, where $\mu,\nu$ takes values in $\{0,x,y,z\}$. For the correlations that vanish, we assume that the corresponding decay exponent $\alpha$ is infinite. Thus, we can write the correlation function as:
@@
where $c_{\mu\nu}\equiv\text{Tr}(\rho_{1}\sigma_{\mu}^{1})\text{Tr}(\rho_{r}\sigma_{\nu}^{r})$ is independent of $r$ due to translational invariance. Using this, the total nonstabilizerness of the state can be expressed as:
@@
where $(1+x)^{-1}\sim(1-x)$ and $\text{log}(1+x)\sim x$ for small $x$ are used. Denoting $\alpha^{*}\equiv\min_{\mu\nu}\alpha_{\mu\nu}$ and $\overline{\alpha}\equiv\min_{\mu\nu}\{\alpha_{\mu\nu};c_{\mu\nu}\neq 0\}$, it follows that the in the large $r$ limit, the total nonstabilizerness $M$ is governed by:
@@
Thus, the decay exponent of the total nonstabilizerness is given by:
@@
At the critical point of the TFIM, we have $\alpha^{*}=\alpha_{xx}=1/4$ and $\overline{\alpha}=\alpha_{zz}=2$ [7], which results in $\alpha_{M}=1/2$, identical to the decay exponent of the mutual information. Furthermore, since the ground state is in canonical form, the NN is equivalent to the total nonstabilizerness of the two-qubit state. This completes the proof.
@@
In this section, we provide numerical details for the simulations conducted in the main text. The ground state of the TFIM is obtained using the density matrix renormalization group algorithm using the ITensor package [8, 9], with the maximum bond dimension $\chi=500$ and the truncation cutoff $\epsilon=10^{-12}$. For the monitored Haar-random unitary circuit at the critical point, the time evolution was performed by the time-evolving block decimation algorithm, with the maximum bond dimension $\chi=1000$ and the truncation cutoff $\epsilon=10^{-6}$.
When calculating the NN via numerical optimization over all local unitary transformations, we parameterize the local unitary on a single qubit as $U=\begin{pmatrix}\text{cos}(\theta/2)&-e^{i\lambda}\text{sin}(\theta/2)\\ e^{i\phi}\text{sin}(\theta/2)&e^{i(\phi+\lambda)}\text{cos}(\theta/2)\end{pmatrix}$, where $\phi\in[0,2\pi]$, $\theta\in[0,\pi]$ and $\lambda\in[0,2\pi]$. The optimization process employed the Nelder-Mead algorithm, as implemented in the Optim package [10]. To mitigate the risk of converging to a local minimum, we initiated the optimization from 100 distinct starting points and selected the minimal value achieved.
S2.2 Separable mixed states host nonzero NN
In this section, we demonstrate that for any choice of nonstabilizerness measure for mixed states, certain separable states will always exhibit a nonzero NN. A separable state is defined as the convex hull of pure product states [11]:
@@
where $\psi_{i}$ are pure stabilizer states. When STAB is regarded as the set of free states, it specifically captures distillable non-Clifford resources. On the other hand, regarding $\text{STAB}_{0}$ as free states, it also incorporates the non-flat probability distributions of mixed states as resources. In the second SRE used in the main text, $\text{STAB}_{0}$ is treated as the set of free states, while other nonstabilizerness measures, such as the robustness of magic (RoM) [14, 15, 16] and mana [17], consider STAB as the set of free states. It is also worth mentioning that the SRE has been extended by a convex-roof construction to provide an alternative measure in the latter case [18].
@@
This scenario changes when STAB is considered as the set of free states, since any single-qubit density matrix can always be diagonalized, thus ensuring that it lies within STAB. However, we are unaware of any proof guaranteeing that every separable multi-qubit state can be transformed into a state belonging to STAB using only local unitary transformations, and we conjecture that this is not the case. To provide numerical evidence, consider the separable state:
@@
Here, $\left|\phi_{0}\right\rangle=\text{cos}(\pi/8)\left|0\right\rangle+\text{sin}(\pi/8)\left|1\right\rangle$. Using RoM as the nonstabilizerness measure, which is defined as $R(\rho)=\min_{x}\{\sum_{i}|x_{i}|;\rho=\sum_{i}x_{i}\left|\psi_{i}\right\rangle\left\langle\psi_{i}\right|\}-1$, we numerically minimize the nonstabilizerness over all possible local unitary transformations and find that $\mathcal{M}_{R}(\rho_{0})\approx 0.0703$, which is nonzero. This indicates that even when using nonstabilizerness measures that treat STAB as the set of free states, certain separable states can still exhibit a nonzero NN.